Dear all,
please find below a short summary of the activities of the th related to
the AO system.


People at LBTO:


R. Briguglio


remote support
M.Xompero, A. Puglisi (VNC+Skype)



Summary
Because on unfavourable environmental conditions, the dome has not been opened all night long. Software modifications have been implemented and verified on the system:
* new DSP code has been installed and verified;
* automatic dump of masterdiagnostic process was implemented and tested, to keep trace of system faults;
* system GUIs have been modified to avoid sudden deaths (experienced during the last days);
* modification on system processes done to automatically handle the diagnostic variables;
* software modifications to fix bugs on ADC and rerotator.

At the end of the software tests, the shell has been set as a final verification: the operation was successfully. The shell gap at its rest position is on its nominal range (i.e.24um-39um), the "ear" on shell border is not visible.


ciao

Runa

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback